OpenStudy (anonymous):

A man cleaning his apartment pushes a vac- uum cleaner with a force of magnitude 83.6 N. The forcemakes an angle of 17.9 with the hor- izontal floor. The vacuum cleaner is pushed 2.65 m to the right along the floor. Calculate the work done by the 83.6 N force. Answer in units of J

OpenStudy (imtiaz7):

Using formula to find work done: W=F.S As the force acts horizontally then we use cos: W=FS*Cos W=83.6*2.65*Cos17.9 W=221.54*0.9515 W=210.79J

OpenStudy (anonymous):

Thank you :)

OpenStudy (imtiaz7):

welcome @RavenLynette